OpenStudy (anonymous):

how to solve this equation radical sign n + 2 = 2

OpenStudy (anonymous):

to get rid of radical, square both sides, so: sqrt(n+2)^2 = 2^2 n+2 =+/- 4 (this is plus or minus 4, so there's two answers) n=2, and n = -6

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you

OpenStudy (radar):

How does squaring 2 on the right side give you a -4??? I realize -6 would result in a complex number not a integer 2.

OpenStudy (anonymous):

bahaha, i did that backwards. its when you take the radical you have +/-

OpenStudy (radar):

Jenny02, I would double check that solution.\[\sqrt{-6 + 2}\neq 2\]

OpenStudy (anonymous):

good call radar :) no matter how many times i take a root or square something, i still get that backwards. radar is correct.

OpenStudy (radar):

I surely understand. Working with radicals and exponents can be confusing some time. I am glad that you also agree.....it gives me more confidence lol